OpenStudy (jamesreed30):

Find the exact area of a circle with the given circumference 3 pi

OpenStudy (will.h):

Oh so we have thw circumference And in order to find the area we will need to find the radius So Circumstance = 2 * pi * r We have the Circumstance equal to 3pi so 3pi = 2 * pi * r Can you solve for r We need ro find the radius So that we can use it in the formula Area = pi * r^2

OpenStudy (jamesreed30):

Would the answer be 12.56

OpenStudy (sshayer):

i think you are not correct. \[2 \pi r=3 \pi,r=\frac{ 3 \pi }{ 2 \pi }=\frac{ 3 }{ 2 },r^2=\left( \frac{ 3 }{ 2 } \right)^2=\frac{ 9 }{ 4 }\] \[area=\pi r^2=\frac{ 22 }{ 7 }\times \frac{ 9 }{ 4 }=\frac{ 99 }{ 14 }=....\]

OpenStudy (will.h):

How much did you get for the radius

OpenStudy (jamesreed30):

4

OpenStudy (will.h):

That's wrong

OpenStudy (jamesreed30):

What should it be